# RankAlign: A Ranking View of the Generator-Validator Gap in Large Language Models

### v1

@@ C.1 In-domain results
Table 7: Detailed performance metrics across methods for Hypernymy and SWORDS tasks with the Llama-3.2-3B model.
@@
Table 8: Performance metrics across methods for LAMBADA and TriviaQA with Llama-3.2-3B.
@@
Table 9: Detailed performance metrics across methods for SWORDS task with the Llama-3.2-3B-Instruct model.
@@ C.2 Comparison with DPO
Nevertheless, to compare directly with DPO, we explore whether adding comparison with reference model $p_{ref}$ would bring performance gains in out setting. Specifically, we rewrite the loss functions as follows:
@@
We report the results of RankAlign and RankAlign-V2G with a reference term in the loss function in Tables 11 and 12. We note that the overall effect of using a reference term in the loss function is neutral to negative.
@@ C.3 Cross-domain results
Ranking G2V 80.9 73.1 57.9 91.5 85.1 21.7 1.6
Ranking V2G 76.5 54.9 43.2 96.8 83.6 28.0 1.7
@@
Ranking G2V 85.0 64.9 65.7 96.1 85.3 9.4 0.9
Ranking V2G 80.7 57.1 49.5 96.0 85.4 14.1 1.0
[…]
@@ C.4 G-V Consistency results
We evaluate G-V Consistency according to Li et al. (2024b). Specifically, we first prompt the model with generator prompts $x_{G}$ and set the corresponding $\hat{y_{A}}$ as the generator output. Then we calculate G-V Consistency as $\mathbb{E}[\mathbb{1}[y_{V}=\text{Yes}]]$, where $y_{V}\sim p_{LM}(\cdot\mid T(x_{G},\hat{y_{A}}))$. We compare model performance in terms of Log-odds Correlation and G-V Consistency between the base model and Consistency FT-ed models.
@@ Appendix D Prompt Variations
Table 16: Generator and validator prompt variants and their associated cross-prompt correlations for the Hypernymy task. Here RankAlign was trained on the generator-validator prompt templates $G(z):=$“Complete the sentence: $z$ are a kind of”, and $V(z,y_{A}):=$“Do you think $z$ are $y_{A}$? Answer:”
@@ Appendix E Prompt templates for experiments
We present prompt templates for each task in Table 171819, and 21 respectively. Optional exemplars in validator prompts are shown to non-instruct-tuned models.

### v2

@@ C.1 In-domain results
RankAlign-$\alpha$ 84.1 70.4 70.6 90.5 78.5 27.0 3.5
Table 7: Performance metrics across methods for Hypernymy and SWORDS tasks with the Llama-3.2-3B model.
@@
RankAlign-$\alpha$ 70.8 13.3 85.5 40.0
Table 8: Performance metrics across methods for LAMBADA and TriviaQA with the Llama-3.2-3B model
Llama-3.2-3B-Instruct
@@
Llama-3.1-8B
The in-domain consistency and accuracy results for experiments on Hypernymy and SWORDS with Llama-3.1-8B are given in Table 11. Results for LAMBADA and TriviaQA for Llama-3.1-8B are in Table 12.
@@
RankAlign-$\alpha$ 72.1 57.5 49.6 90.1 76.8 24.1 3.0
Table 9: Performance metrics across methods for the SWORDS task with the Llama-3.2-3B-Instruct model.
@@
Hypernym Base 77.9 58.9 43.6 97.7 83.7 13.2 1.2
SFT 89.3 58.9 61.8 98.1 51.5 73.9 11.7
Consistency FT 79.8 51.7 47.1 97.9 58.9 69.9 12.9
DPO-G2V 81.1 53.2 56.9 97.7 81.7 56.9 3.3
DPO-V2G 76.1 45.2 76.6 82.8 82.9 12.7 1.3
RankAlign 94.0 89.1 88.4 93.8 83.0 12.9 1.2
RankAlign-V2G 94.8 77.5 84.1 97.8 91.7 35.1 0.5
RankAlign-$\alpha$ 95.1 88.5 90.8 94.0 82.9 28.3 2.4
SWORDS Base 71.2 44.5 48.1 89.7 77.6 29.9 3.7
SFT 71.5 40.9 46.1 90.7 74.6 40.1 4.7
Consistency FT 69.8 39.9 45.1 89.5 76.3 38.1 4.5
DPO-G2V 71.0 44.1 46.7 90.0 78.7 30.9 3.6
DPO-V2G 70.2 42.9 41.8 91.6 78.0 30.6 3.8
RankAlign 81.4 69.1 61.3 90.5 77.5 29.0 3.6
RankAlign-V2G 76.5 50.9 58.8 89.7 79.5 30.7 3.3
RankAlign-$\alpha$ 70.2 42.2 44.8 90.1 81.0 26.4 3.4
Table 11: Performance metrics across methods for Hypernymy and SWORDS tasks with the Llama-3.1-8B model.
Task Method $\rho$-pos R@0 A@100 MRR-P
LAMBADA Base 0.6 79.6 99.9 82.5
SFT 6.9 100.0 99.7 86.7
Consistency FT 4.8 100.0 99.8 86.1
DPO-V2G 62.7 99.3 99.9 82.3
RankAlign 65.2 74.9 92.1 61.6
RankAlign-V2G 33.5 38.2 4.1 0.5
RankAlign-$\alpha$ 26.2 30.1 99.9 82.5
TriviaQA Base 32.2 100.0 92.8 62.7
SFT 36.9 100.0 95.4 69.9
Consistency FT 43.7 100.0 94.9 72.2
DPO-V2G 50.1 77.6 92.5 62.7
RankAlign 65.2 0.0 92.1 61.6
RankAlign-V2G 43.4 100.0 90.3 57.3
RankAlign-$\alpha$ 63.6 2.5 92.6 62.7
Table 12: Performance metrics across methods for LAMBADA and TriviaQA with the Llama-3.1-8B model
@@ C.2 Comparison with DPO
Nevertheless, to compare directly with DPO, we explore whether adding comparison with reference model $p_{ref}$ would bring performance gains in out setting. Specifically, we change the loss functions as follows:
@@
We report the results of RankAlign and RankAlign-V2G with a reference term in the loss function in Tables 13 and 14. We note that the overall effect of using a reference term in the loss function is neutral to negative.
@@ C.3 Cross-domain results
RankAlign 80.9 73.1 57.9 91.5 85.1 21.7 1.6
RankAlign-V2G 76.5 54.9 43.2 96.8 83.6 28.0 1.7
RankAlign-$\alpha$ 84.4 73.9 66.4 94.1 82.1 36.5 2.0
@@
RankAlign 85.0 64.9 65.7 96.1 85.3 9.4 0.9
RankAlign-V2G 80.7 57.1 49.5 96.0 85.4 14.1 1.0
RankAlign-$\alpha$ 86.9 69.5 67.9 96.5 85.5 17.5 1.2
Table 15: Cross-domain results, $\text{SWORDS}\rightarrow\text{Hypernymy}$, for Gemma-2-2B and Llama-3.2-3B.
@@
Results on the generalization of RankAlign-$\alpha$ across the four tasks are in Table 16 and show a similar trend to the RankAlign results in Table 3.
Hypernymy Gemma-2 12.9 27.3 38.0 16.5 24.7 16.6 -12.9 13.7
Llama-3.2 16.8 33.9 46.1 13.6 15.2 13.7 -4.4 25.8
SWORDS Gemma-2 8.1 19.0 21.3 22.6 36.4 30.6 2.9 26.0
Llama-3.2 8.4 13.8 23.8 30.3 43.4 38.1 -6.0 33.3
LAMBADA Gemma-2 -7.5 -4.2 -9.5 -44.9 -24.4 -42.9 51.1 -21.2
Llama-3.2 1.5 1.3 1.6 -3.3 7.3 -4.5 27.8 8.9
TriviaQA Gemma-2 3.7 6.5 5.2 -1.3 5.1 -3.5 -5.7 53.7
Llama-3.2 2.2 4.3 9.8 16.8 24.7 17.0 -13.8 57.8
Table 16: Cross-dataset evaluation showing the difference in $\rho$ scores (between the RankAlign-$\alpha$ trained model and the base LM) when training on a dataset (row) and evaluating on target datasets (column). Scores greater than 0 indicate that RankAlign-$\alpha$ generalizes across tasks. Values in gray are the in-domain results.
@@ C.4 G-V Consistency results
We evaluate G-V Consistency according to Li et al. (2024b). Specifically, we first prompt the model with generator prompts $x_{G}$ and set the corresponding $\hat{y_{A}}$ as the generator output. Then we calculate G-V Consistency as $\mathbb{E}[\mathbb{1}[y_{V}=\text{Yes}]]$, where $)y_{V}\sim p_{LM}(\cdot\mid T(x_{G},\hat{y_{A}}$)). We compare model performance in terms of Log-odds Correlation and G-V Consistency between the base model and Consistency FT-ed models.
@@ Appendix D Prompt Variations
Table 19: Generator and validator prompt variants and their associated cross-prompt correlations for the Hypernymy task. Here RankAlign was trained on the generator-validator prompt templates G(z):=G(z):=“Complete the sentence: $z$ are a kind of”, and $V(z,y_{A}):=$“Do you think $z$ are $y_{A}$? Answer:”
@@ Appendix E Prompt templates for experiments
We present prompt templates for each task in Table 202122, and 24 respectively. Optional exemplars in validator prompts are shown to non-instruct-tuned models.